The weight of a synthetic ball varies directly with the cube if its radius. A ball with a radius of 2 inches weighs 440 pounds. Find the weight of a ball witha 3 inch radius?
Answer by Alan3354(69443)

440 = k*2^3 = 8k
k = 55
------------
W = k*3^3 = 27k
W = 1485 pounds